OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x)

OpenStudy (aakashsudhakar):

f-1(x) represents the inverse function of f(x). The easiest way to do this is through a little cheap trick: switch the f(x) and x positions on the function and then isolate f(x). You're given that f(x) = x^2 - 16 Switch f(x) and x. x = f(x)^2 - 16 Now isolate f(x). f(x)^2 = x + 16 f(x) = ±√(x + 16) Now simply erase f(x) and rewrite it as f-1(x), since this is now the inverse function of f(x). f-1(x) = ±√(x + 16) NOTE: Don't forget, when taking the square root, always remember to put ± before the expression!

OpenStudy (anonymous):

f^(-1) (x) is the inverse function of f. To find the inverse function, interchange x with y, and y with x, AND then solve for y. so the inverse will be x = y^2 - 16 now we solve for y, we get y^2 = x + 16 y = + or - sqrt(x+16) NOTE: The inverse y = +/- sqrt(x+16) is NOT a function. The inverse of a function NEED NOT be a function, as demonstated in this exercise.